OpenStudy (anonymous):

how to solve this system of equations with substitution? y= 4x x + y = 5

OpenStudy (lin.ivory):

X+4X=5 5X=5 5/5=1 X=1

OpenStudy (anonymous):

y = 4x is equation1. x + y = 5 can also be written to make y the subject. y = 5 - x is equation 2 EQUATION 1 --> y = 4x EQUATION 2 --> y = 5 - x This basically means: 4x = 5 - x 4x + x = 5 5x = 5 x = 1